Prove that log2 3 is irrational.
1
Expert's answer
2020-11-27T11:46:02-0500

Suppose that "\log_23" is a rational number. Since "\log_23>0," "\log_23=\frac{a}{b},"

where "a,b\in\mathbb N." Then "3=2^{\frac{a}{b}}," and therefore, "3^b=2^a." But the last equality is impossible beacause the natural number "2^a" is divisible by 2, but the number "3^b" is not divisible by 2. This contradiction proves that the assumption was wrong, hence "\log_23" is irrational.